# ParamANN: A Neural Network to Estimate Cosmological Parameters for $\Lambda$CDM Universe Using Hubble Measurements

### v2

@@ Abstract
In this article, we employ a machine learning (ML) approach for the estimations of four fundamental parameters, namely, the Hubble constant ($H_{0}$), matter ($\Omega_{0m}$), curvature ($\Omega_{0k}$) and vacuum ($\Omega_{0\Lambda}$) densities of non-flat $\Lambda$CDM model. We use $31$ Hubble parameter values measured by differential ages (DA) technique in the redshift interval $0.07\leq z\leq 1.965$. We create an artificial neural network (ParamANN) and train it with simulated values of $H(z)$ using various sets of $H_{0}$, $\Omega_{0m}$, $\Omega_{0k}$, $\Omega_{0\Lambda}$ parameters chosen from different and sufficiently wide prior intervals. We use a correlated noise model in the analysis. We demonstrate accurate validation and prediction by ParamANN. ParamANN provides an excellent cross-check for the validity of the $\Lambda$CDM model. We obtain $H_{0}=68.14\pm 3.96$ $\rm{kmMpc^{-1}s^{-1}}$, $\Omega_{0m}=0.3029\pm 0.1118$, $\Omega_{0k}=0.0708\pm 0.2527$ and $\Omega_{0\Lambda}=0.6258\pm 0.1689$ by using the trained network. These parameter values agree very well with the results of global CMB observations of the Planck collaboration. We compare the cosmological parameter values predicted by ParamANN with those obtained by the MCMC method. Both the results agree well with each other. This demonstrates that ParamANN is an alternative and complementary approach to the well-known Metropolis-Hastings algorithm for estimating the cosmological parameters by using Hubble measurements.
aDepartment of Physics, Indian Institute of Science Education and Research Bhopal,
Bhopal - 462066, Madhya Pradesh, India
Keywords: Hubble parameter - Cosmological density parameters - Machine learning - MCMC
@@ 1 Introduction
Measurements of Hubble expansion rate, $H(z)$, at different redshifts ($z$), provide a powerful probe to study the evolution of the Universe due to the relatively simple relationship between $H(z)$ and the cosmological parameters. This has the potential to lead to easier interpretation of the results of an analysis by unambiguously identifying the underlying cosmological model. Although a simple 6 parameter $\Lambda$ cold dark matter ($\Lambda$CDM) model is generally believed to outweigh other cosmological models, there are some indications that $\Lambda$CDM model may not be consistent with respect to all different observational probes. The area of incompatibility is known as the so-called problem of ‘Hubble tension’. The tension was reported since $H_{0}$ values estimated by local and global observations do not agree with each other. The value of $H_{0}$ constrained by Planck collaboration [2] shows $\sim 3.6\sigma$ tension with the same measured by local observational data, in which local measurements of $H_{0}$ demand the higher value of this parameter. Recent local observations of Hubble Space Telescope(HST) [3, 4, 5, 6, 7] estimate the value of the Hubble constant which shows approximately $4$-$5\sigma$ tension with Planck’s estimation of $H_{0}$. Local measurement of $H_{0}$ by using gravitationally lensed quasars, i.e. H0LiCOW [8], shows $\sim 5.3\sigma$ tension with Planck’s $H_{0}$ value. Di Valentino (2021) [9] estimated the value of $H_{0}$ by combining 23 local measurements of this parameter and this estimation shows $5.9\sigma$ tension with $H_{0}$ constrained by CMB observations [2]. An alternative method to estimate $H_{0}$ value (along with other cosmological parameters) is of utmost importance in contemporary cosmological analysis.
@@
It is worth mentioning that the problem of Hubble tension seems to be somewhat perplexing in nature. On one hand, some observations [10] indicate a high value of $H_{0}$. Riess et al. (2022) [7] constrain $H_{0}=73.04$ $\rm{kmMpc^{-1}s^{-1}}$ with a uncertainty of 1.04 $\rm{kmMpc^{-1}s^{-1}}$ using Cepheid-SNe samples for their baseline redshift range $0.150.0233<z<$ 0.15. The Pantheon+ analysis [11] use 1550 distinct type Ia Supernovae in the redshift interval 0.001 to 2.26 and report the $H_{0}=73.5\pm 1.1$ $\rm{kmMpc^{-1}s^{-1}}$. Thus it appears that the Hubble tension extends up to a redshift of 2.26 if Supernovae data is used. Interestingly, using the red giant branch (TRGB) calibration to a sample of Type Ia Supernovae, Freedman (2021) [12] reports $H_{0}=69.8\pm 0.6$ (stat) ±$\pm 1.6$ (sys) $\rm{kmMpc^{-1}s^{-1}}$ which is consistent with the $H_{0}$ value estimated by Planck collaboration [2]. Kelly et al. (2023) [13] obtain $H_{0}=64.8^{+4.4}_{-4.3}$ $\rm{kmMpc^{-1}s^{-1}}$ by using eight lens models and $66.6^{+4.1}_{-3.3}$ $\rm{kmMpc^{-1}s^{-1}}$ from two preferred models. A recent article [14] report $H_{0}=67.6^{+4.3}_{-4.2}$ $\rm{kmMpc^{-1}s^{-1}}$ using VLBI and gravitational wave observations from bright binary black hole GW190521 and conclude that the value is consistent with the Planck observations, relaxing the tension. At the backdrop of such diverse indications from the observations, it is a very important task to measure the cosmological parameters in an as much general framework (i.e., measuring all the independent cosmological parameters) as possible using the favoured $\Lambda$CDM model and new observational data. Leaf & Melia (2017) [15] analyze only differential ages (DA) Hubble measurements using a two-point diagnostic for model comparison since the measurements of $H(z)$ from cosmic chronometers are model independent. Geng et al. (2018) [16] use $H(z)$ data measured by DA and baryon acoustic oscillation (BAO) techniques to constrain the cosmological parameters as well as quantify the impact of future $H(z)$ measurements in the estimation of these parameters. The recent literatures [17, 18, 19, 20, 21, 22] effectively use the low redshift data (i.e., DA+BAO Hubble data, QSO angular size, Pantheon, DES supernova etc.) to analyse the various cosmological models. In this article, we use $H(z)$ observations alone to constrain the parameters using artificial intelligence (AI) as the driving engine.
@@
We use AI to estimate the Hubble constant ($H_{0}$) and today’s density parameters ($\Omega_{0m}$, $\Omega_{0k}$ and $\Omega_{0\Lambda}$) for $\Lambda$CDM Universe from the $H(z)$ values measured by DA technique [23]. We create an ANN (hereafter ParamANN) for modelling a direct mapping function between the observed $H(z)$ and corresponding four parameters of the $\Lambda$CDM Universe. The density parameters estimated by ParamANN indicate the spatially flat $\Lambda$CDM Universe which is consistent with the results of Planck collaboration [2]. Moreover, the Hubble constant predicted by ParamANN agrees excellently with Planck’s estimation of the same. Furthermore, we compare the parameter values predicted by ParamANN with the same estimated by the Markov chain Monte Carlo (MCMC) approach, e.g. Metropolis-Hastings algorithm [24]. We notice that both approaches provide similar values with similar levels of uncertainty for each parameter. Therefore, ParamANN can be used as an alternative to the traditional approach for estimating the fundamental parameters from Hubble data. We note that whether the MCMC analysis depends on the specific likelihood function, our efficient and simple neural network (for estimating cosmological parameters) performs as a likelihood-free inference method [25].
@@
The primary motivations of our current article stem from the perspective of both theoretical and observational fronts. From the observational side, the problem like Hubble tension [2, 3, 4, 5, 6, 7, 8, 9, 10, 11] exists and needs further understanding using various types of available data. AI has become one of the promising tools to investigate the observed data since an ML model can predict (in principle) a complicated function once it has been trained successfully. Moreover, there are several articles in the literature [26, 27, 28, 29, 30, 31, 32, 15, 16, 18, 33, 34, 35, 36, 37] which try to measure the cosmological parameters assuming flat spatial curvature of our Universe. However, it is also important to ask whether we can estimate these cosmological parameters by relaxing the assumption of the flatness of the Universe. In this article, we motivate ourselves to estimate cosmological density parameters along with today’s Hubble parameter value for a general non-flat $\Lambda$CDM Universe. From the observational perspective, many new experiments are being proposed, e.g., Echo (aka CMB-Bharat [^3] http://cmb-bharat.in/), CCAT-prime [38], PICO [39], Lite-Bird [40], SKA [41], which has lower noise level implying the model parameters of a theory can be measured with higher accuracy. The higher accuracy of the future generation observations demands accurate constraining of cosmological parameters to distinguish between different models of the Universe which is a major step to finding a better and more accurate theoretical understanding of the physics of the Universe.
@@
In the modern era, ML techniques are utilized as the powerful tools to analyze the observational data of the several fields in cosmology [42, 43]. Instead of Metropolis-Hastings algorithm, artificial neural network (ANN) can be used as an alternative approach for the Bayesian inference in cosmology effectively reducing the computational time [44, 45, 46, 47]. Moreover, ANNs can be performed for the non-parametric reconstructions of cosmological functions [48, 49, 50, 51]. For the estimations of cosmological parameters by using CMB data, Mancini et al. (2022) [52] implement several types of ANN and show less computational time in the Bayesian process by using these ANNs. Baccigalupi et al. (2000) [53] utilize ANN to separate different types of foregrounds (e.g., thermal dust emissions, galactic synchrotron, and radiation emitted by galaxy clusters) from CMB signal. Petroff et al. (2020) [54] develop a Bayesian spherical convolutional neural network (CNN) to recover the CMB anisotropies from the foreground contaminations. Using a CNN, Shallue & Einstein (2023) [55] reconstruct the cosmological initial conditions from the late-time and non-linearly evolved density fields. ML can be used to reconstruct full sky CMB temperature anisotropies from the partial sky maps [56, 57]. Khan & Saha (2023) [58] develop an ANN to estimate the dipole modulation from the foreground cleaned CMB temperature anisotropies. In our previous article [59], we use a CNN to recover full sky $E$- and $B$-modes polarizations from the partial sky maps avoiding the so-called $E$-to-$B$ leakage. Wang et al. (2020b) [60] use ANN to estimate cosmological parameters from the temperature power spectrum of CMB. Bengaly et al. (2023) [33] employ ML algorithm to constrain the Hubble constant ($H_{0}$) by using DA Hubble measurements. Moreover, the recent literature [34, 35, 36, 37, 61, 62] also perform the ML techniques to analyze the low redshift data (i.e., Hubble measurements (DA+BAO), Type Ia Supernovae etc.) for constraining the cosmological parameters of present Universe.
@@
We organize our article as the following. In section 2, we describe the relation between Hubble parameter and redshift for $\Lambda$CDM Universe. In section 3, we present the observed $H(z)$ data, discuss the signal and noise models used in our analysis, and describe the deep learning of ParamANN respectively. In this section, we also discuss the MCMC method. In section 4, we show the predicted parameter values for the test set as well as for the observed $H(z)$ data. We compare the results predicted by ParamANN with the results obtained by MCMC. Finally, in section 5, we conclude with the beneficial discussions of our analysis.
@@ 2 Formalism
Einstein’s equations which can be encapsulated into a compact form using tensor notations follow
@@
In spherical coordinates, the Friedmann-Robertson-Walker (FRW) line element can be written as
[…]
@@
[…]
where $\rho(t)$ and $P(t)$ denote the density and gravitational pressure of the Universe respectively.
Using equations (3) and (4) as well as the equation of state $P(t)=\omega c^{2}\rho(t)$, we obtain the energy-momentum conservation law which is given by
$\partialderivative{\rho}{t}+3\left(1+\omega\right)H(t)\rho(t)$ $=$ $0,$ (5)
[…]
@@
where $z$ denotes the redshift and ‘0’ stands for present Universe (i.e., $z=0$). In equations (6), (7) and (8), subscript notations $m,r$ and $\Lambda$ define the matter, radiation and vacuum densities respectively. Moreover, redshift is related to scale factor as $(1+z)=a_{0}/a$.
Neglecting the radiation density for late time Universe, equation (3) can be written as
@@
In equation (9), $\Omega$ denotes the density parameter which is defined by $\rho/\rho_{c}$ for matter and vacuum densities, where $\rho_{c}$ is called the critical density which is expressed as $3H^{2}/8\pi G$. However, the curvature density parameter ($\Omega_{k}$) is defined by $-kc^{2}/a^{2}H^{2}$.
@@
Using equation (11), to apply the condition of present Universe on the curvature density parameter, equation (10) can be written as

### v3

@@ Abstract
In this article, we employ a machine learning (ML) approach for the estimations of four fundamental parameters, namely, the Hubble constant ($H_{0}$), matter ($\Omega_{0m}$), curvature ($\Omega_{0k}$) and vacuum ($\Omega_{0\Lambda}$) densities of non-flat $\Lambda$CDM model. We use $31$ Hubble parameter values measured by differential ages (DA) technique in the redshift interval $0.07\leq z\leq 1.965$. We create an artificial neural network (ParamANN) and train it with simulated values of $H(z)$ using various sets of $H_{0}$, $\Omega_{0m}$, $\Omega_{0k}$, $\Omega_{0\Lambda}$ parameters chosen from different and sufficiently wide prior intervals. We use a correlated noise model in the analysis. We demonstrate accurate validation and prediction using ParamANN. ParamANN provides an excellent cross-check for the validity of the $\Lambda$CDM model. We obtain $H_{0}=68.14\pm 3.96$ $\rm{kmMpc^{-1}s^{-1}}$, $\Omega_{0m}=0.3029\pm 0.1118$, $\Omega_{0k}=0.0708\pm 0.2527$ and $\Omega_{0\Lambda}=0.6258\pm 0.1689$ by using the trained network. These parameter values agree very well with the results of global CMB observations of the Planck collaboration. We compare the cosmological parameter values predicted by ParamANN with those obtained by the MCMC method. Both the results agree well with each other. This demonstrates that ParamANN is an alternative and complementary approach to the well-known Metropolis-Hastings algorithm for estimating the cosmological parameters by using Hubble measurements.
  • October 2024
Keywords: Hubble parameter, cosmological density parameters, machine learning, MCMC
@@ 1 Introduction
Measurements of Hubble expansion rate, $H(z)$, across various redshifts ($z$), offer a valuable tool to study the Universe’s evolution due to the relatively simple relationship between the $H(z)$ and fundamental parameters of cosmology. This has the potential to lead to easier interpretation of the results of an analysis by unambiguously identifying the underlying cosmological model. Although a simple 6 parameter $\Lambda$ cold dark matter ($\Lambda$CDM) model is generally believed to outweigh other cosmological models, some observations suggest potential discrepancies between the $\Lambda$CDM model and certain experimental data. The area of incompatibility is known as the so-called problem of ‘Hubble tension’. The tension was reported since $H_{0}$ values estimated by local and global observations do not agree with each other. The value of $H_{0}$ constrained by Planck collaboration [1] shows $\sim 3.6\sigma$ tension with the same measured by local observational data, in which local measurements of $H_{0}$ demand the higher value of this parameter. Recent local observations of Hubble Space Telescope(HST) [2, 3, 4, 5, 6] estimate the value of $H_{0}$ which shows approximately $4$-$5\sigma$ tension with Planck’s estimation of $H_{0}$. Local measurement of $H_{0}$ by using gravitationally lensed quasars, i.e. H0LiCOW [7], shows $\sim 5.3\sigma$ tension with Planck’s $H_{0}$ value. Di Valentino (2021) [8] estimated the value of $H_{0}$ by combining 23 local measurements of this parameter and this estimation shows $5.9\sigma$ tension with $H_{0}$ constrained by CMB observations [1]. An alternative method to estimate $H_{0}$ value (along with other cosmological parameters) is of utmost importance in contemporary cosmological analysis.
@@
It is worth mentioning that the problem of Hubble tension seems to be somewhat perplexing in nature. On one hand, some observations [9] indicate a high value of $H_{0}$. Riess et al(2022) [6] constrain $H_{0}=73.04$ $\rm{kmMpc^{-1}s^{-1}}$ with a uncertainty of 1.04 $\rm{kmMpc^{-1}s^{-1}}$ using Cepheid-SNe samples for their baseline redshift range $0.150.0233<z<$ 0.15. The Pantheon+ analysis [10] use 1550 distinct type Ia Supernovae in the redshift interval 0.001 to 2.26 and report the $H_{0}=73.5\pm 1.1$ $\rm{kmMpc^{-1}s^{-1}}$. Thus it appears that the Hubble tension extends up to a redshift of 2.26 if Supernovae data is used. Interestingly, using the red giant branch (TRGB) calibration to a sample of Type Ia Supernovae, Freedman (2021) [11] reports $H_{0}=69.8\pm 0.6$ (stat) ±$\pm 1.6$ (sys) $\rm{kmMpc^{-1}s^{-1}}$ which is consistent with the $H_{0}$ value estimated by Planck collaboration [1]. Kelly et al(2023) [12] obtain $H_{0}=64.8^{+4.4}_{-4.3}$ $\rm{kmMpc^{-1}s^{-1}}$ by using eight lens models and $66.6^{+4.1}_{-3.3}$ $\rm{kmMpc^{-1}s^{-1}}$ from two preferred models. A recent article [13] report $H_{0}=67.6^{+4.3}_{-4.2}$ $\rm{kmMpc^{-1}s^{-1}}$ using VLBI and gravitational wave observations from bright binary black hole GW190521 and conclude that the value is consistent with the Planck observations, relaxing the tension. At the backdrop of such diverse indications from the observations, it is a very important task to measure the cosmological parameters in an as much general framework (i.e., measuring all the independent cosmological parameters) as possible using the favored $\Lambda$CDM model and new observational data. Leaf & Melia (2017) [14] analyze only differential ages (DA) Hubble measurements using a two-point diagnostic for model comparison since the measurements of $H(z)$ from cosmic chronometers are model independent. Geng et al(2018) [15] use $H(z)$ data measured by DA and baryon acoustic oscillation (BAO) techniques to constrain the cosmological parameters as well as quantify the impact of future $H(z)$ measurements in the estimation of these parameters. The recent literatures [16, 17, 18, 19, 20, 21] effectively use the low redshift data (i.e., DA+BAO Hubble data, QSO angular size, Pantheon, DES supernova etc.) to analyze the various cosmological models. Almeida et al(2017) [22] investigated the inflationary phase and other stages of the Universe within a redshift-dependent Lorentz-violating time-like background and noticed that the dark energy era at low redshifts is dominated by this background. Aktaş et al(2012) [23] demonstrated that within the framework of $f(R)$ gravity, dark energy, modeled as both a perfect fluid and a mesonic scalar field in Marder space-time, exhibits phantom-like behavior with an equation of state $\omega<-1$. Iqbal & Fawad (2019) [24] examined Tsallis, Renyi, and Sharma–Mittal holographic dark energy models in a flat Friedmann-Robertson-Walker (FRW) universe, incorporating interactions between dark energy and cold dark matter under the DGP braneworld framework. They have shown that these models agree with the accelerated expansion and stability, consistent with observational data and the $\Lambda$CDM model. In this article, we use $H(z)$ observations alone to constrain the parameters of $\Lambda$CDM model using artificial intelligence (AI) as the driving engine.
@@
We use AI to estimate the Hubble constant ($H_{0}$) and today’s density parameters ($\Omega_{0m}$, $\Omega_{0k}$ and $\Omega_{0\Lambda}$) for $\Lambda$CDM Universe from the $H(z)$ values measured by DA technique [25]. We create an ANN (hereafter ParamANN) for modeling a direct mapping function between the observed $H(z)$ and the corresponding four parameters of the $\Lambda$CDM Universe. The density parameters estimated by ParamANN indicate the spatially flat $\Lambda$CDM Universe which is consistent with the results of Planck collaboration [1]. Moreover, the Hubble constant predicted by ParamANN agrees excellently with Planck’s estimation of the same. Furthermore, we compare the parameter values predicted by ParamANN with the same estimated by the Markov chain Monte Carlo (MCMC) approach, e.g. Metropolis-Hastings algorithm [26]. We notice that both approaches provide similar values with similar levels of uncertainty for each parameter. Therefore, ParamANN can be used as an alternative to the traditional approach for estimating the fundamental parameters from Hubble data. We note that whether the MCMC analysis depends on the specific likelihood function, our efficient and simple neural network (for estimating cosmological parameters) performs as a likelihood-free inference method [27].
@@
The primary motivations of our current article stem from the perspective of both theoretical and observational fronts. From the observational side, the problem like Hubble tension [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] exists and needs further understanding using various types of available data. AI has become one of the promising tools for investigating the observed data since an ML model can predict (in principle) a complicated function once it has been trained successfully. Moreover, there are several articles in the literature [28, 29, 30, 31, 32, 33, 34, 14, 15, 17, 35, 36, 37, 38, 39] which try to measure the cosmological parameters assuming flat spatial curvature of our Universe. However, it is also important to ask whether we can estimate these cosmological parameters by relaxing the assumption of the flatness of the Universe. In this article, we motivate ourselves to estimate cosmological density parameters along with today’s Hubble parameter value for a general non-flat $\Lambda$CDM Universe. From the observational perspective, many new experiments are being proposed, e.g., Echo (aka CMB-Bharat [^2] http://cmb-bharat.in/), CCAT-prime [40], PICO [41], Lite-Bird [42], SKA [43], which has lower noise level implying the model parameters of a theory can be measured with higher accuracy. The higher accuracy of the future generation observations demands accurate constraining of cosmological parameters to distinguish between different models of the Universe which is a major step to finding a better and more accurate theoretical understanding of the physics of the Universe.
@@
In the modern era, ML techniques are utilized as powerful tools to analyze the observational data of several fields in cosmology [44, 45]. Instead of the Metropolis-Hastings algorithm, an artificial neural network (ANN) can be used as an alternative approach for the Bayesian inference in cosmology effectively reducing the computational time [46, 47, 48, 49]. Moreover, ANNs can be performed for the model-free reconstructions of the functions related to cosmology [50, 51, 52, 53]. For the parameter estimation of cosmology by using CMB data, Mancini et al(2022) [54] implement several types of ANN and show less computational time in the Bayesian process by using these ANNs. Baccigalupi et al(2000) [55] utilize ANN to separate different types of foregrounds (e.g., thermal dust emissions, galactic synchrotron, and radiation emitted by galaxy clusters) from CMB signal. Petroff et al(2020) [56] develop a Bayesian spherical convolutional neural network (CNN) to recover the CMB anisotropies from the foreground contaminations. Using a CNN, Shallue & Einstein (2023) [57] reconstruct the initial conditions of the Universe from the late-time density fields which evolve non-linearly. ML can be employed to reconstruct full sky CMB temperature anisotropies from the partial sky maps [58, 59]. Khan & Saha (2023) [60] develop an ANN to estimate the dipole modulation from the foreground cleaned CMB temperature anisotropies. In our previous article [61], we use a CNN to recover full sky $E$- and $B$-modes polarizations from the partial sky maps avoiding the so-called $E$-to-$B$ leakage. Wang et al(2020b) [62] use ANN to estimate cosmological parameters from the temperature power spectrum of CMB. Bengaly et al(2023) [35] employ an ML algorithm to constrain the Hubble constant ($H_{0}$) by using DA Hubble measurements. Moreover, the recent literature [36, 37, 38, 39, 63, 64] also perform the ML techniques to analyze the low redshift data (i.e., Hubble measurements (DA+BAO), Type Ia Supernovae etc.) for constraining the cosmological parameters of present Universe. The literature [65] explored how memristive neural networks (MNNs) with leakage and mixed delays achieve finite-time synchronization (FTS) utilizing the state feedback and adaptive control strategies to establish adequate conditions based on the Filippov solution and the Lyapunov functional technique. Recent studies have demonstrated the potential of machine learning in cosmology. Escamilla-Rivera et al(2020) [50] introduced RNN+BNN, combining recurrent and Bayesian neural networks to study dark energy models using supernovae data, reducing computational load while accounting for uncertainties. Tilaver et al(2021) [66] applied LSTM and the Fisher Information Matrix to predict the Hubble parameter within the Chaplygin gas framework, and Salti et al(2021) [67] investigated the progression of the temperature of the CMB radiation by applying Deep Neural Networks for the modified generalized Chaplygin gas model. The recent work [68] utilized deep learning with a tuned LSTM to predict Hubble parameters from cosmic chronometers. In this paper, we develop ParamANN, an ANN for directly modeling the mapping between $H(z)$ and four parameters of $\Lambda$CDM Universe.
@@
We organize our article as the following. In section 2, we describe the relation between the Hubble parameter and redshift for $\Lambda$CDM Universe. In section 3, we present the observed $H(z)$ data, discuss the signal and noise models used in our analysis, and describe the deep learning of ParamANN respectively. In this section, we also discuss the MCMC method. In section 4, we show the predicted parameter values for the test set as well as for the observed $H(z)$ data. We compare the results predicted by ParamANN with the results obtained by MCMC. Finally, in section 5, we conclude with the beneficial discussions of our analysis.
@@ 2 Formalism
Einstein’s equations, which can be encapsulated into a compact form using tensor notations, are as follows
@@
In spherical coordinates, the FRW line element can be written as
$\textrm{d}s^{2}$ $=$ $c^{2}\textrm{d}t^{2}-a^{2}(t)\left[\frac{\textrm{d}r^{2}}{1-kr^{2}}+r^{2}\textrm{d}l^{2}\right],$ (2)
where $r$, $a(t)$ and $k$ define the radial comoving coordinate, cosmological scale factor and curvature constant of the Universe respectively. The positive, negative, and zero values of $k$ indicate spatially closed, open, and flat Universes respectively. Moreover, $\rm{d}l^{2}$ is defined as
$\textrm{d}l^{2}$ $=$ $\textrm{d}r^{2}+\sin^{2}\theta\textrm{d}\phi^{2},$ (3)
where $\theta$ and $\phi$ are the angular comoving coordinates of the spherical Universe.
@@
$\frac{1}{a^{2}(t)}\left[\left(\frac{\textrm{d}a}{\textrm{d}t}\right)^{2}+kc^{2}\right]$ $=\frac{8\pi G}{3}\rho(t),$ (4)
$\frac{2}{a(t)}\frac{\textrm{d}^{2}a}{\textrm{d}t^{2}}+\frac{1}{a^{2}(t)}\left[\left(\frac{\textrm{d}a}{\textrm{d}t}\right)^{2}+kc^{2}\right]$ $=-\frac{8\pi G}{c^{2}}P(t),$ (5)
where $\rho(t)$ and $P(t)$ denote the density and gravitational pressure of the Universe respectively. Additionally, the equation of state can be written as
$P(t)$ $=$ $\omega c^{2}\rho(t),$ (6)
where $\omega$ denotes the equation of state parameter. Using equations (4), (5) and (6), we obtain the energy-momentum conservation law which is given by
$\frac{\partial\rho}{\partial t}+3\left(1+\omega\right)H(t)\rho(t)$ $=$ $0,$ (7)
where $H(t)$ represents the Hubble parameter which is defined by
$H(t)$ $=$ $\frac{1}{a(t)}\frac{\textrm{d}a}{\textrm{d}t}.$ (8)
The values of $\omega$ are represented as the following
$\omega$ $=$ $\begin{cases}\hskip 13.0pt0&\qquad\text{Matter density},\\ \hskip 13.0pt\frac{1}{3}&\qquad\text{Radiation density},\\ \ -1&\qquad\text{Vacuum density}.\end{cases}$ (9)
Using equation (7), (8) and (9), density components of the Universe are expressed by
@@
where $z$ denotes the redshift and ‘0’ stands for the present Universe (i.e., $z=0$). In equations (10), (11) and (12), subscript notations $m,r$ and $\Lambda$ define the matter, radiation and vacuum densities respectively. Moreover, redshift is related to the scale factor as follows
$(1+z)$ $=$ $\frac{a_{0}}{a},$ (13)
where $a_{0}$ denotes today’s scale factor of the Universe.
Neglecting radiation density for late time Universe, equation (4) can be written as
@@
In equation (14), $\Omega_{m}$ and $\Omega_{\Lambda}$ denote the matter and vacuum density parameters respectively, which are defined by
$\Omega_{m}$ $=$ $\frac{\rho_{m}}{\rho_{c}},$ (15)
$\Omega_{\Lambda}$ $=$ $\frac{\rho_{\Lambda}}{\rho_{c}},$ (16)
where $\rho_{c}$ is called the critical density which is expressed as
$\rho_{c}$ $=$ $\frac{3H^{2}}{8\pi G}.$ (17)
Moreover, the curvature density parameter ($\Omega_{k}$) is defined by
$\Omega_{k}$ $=$ $-\frac{kc^{2}}{a^{2}H^{2}}.$ (18)
@@
Using equation (20), to apply the condition of the present Universe on the curvature density parameter, equation (19) can be written as